OpenStudy (anonymous):

In a certain clothing store, 3 shirts and 5 ties cost $60, and 2 shirts and 3 ties cost $39. What is the cost of each shirt? If s = the cost of a shirt and t = the cost of a tie, which system of equations represents the problem?

OpenStudy (anonymous):

Create tow equations: s+t=60 s+t=39 Now place the number bought respectively with each variable: 3s+5t=60 2s+3t=39 Do you know it from here or do you want me to continue explaining?

OpenStudy (anonymous):

*two

OpenStudy (anonymous):

can you please continue?